# Dimensional Regularization and Two–Loop Vacuum Polarization Operator: Master Integrals, Analytic Results and Energy Shifts

### v1

@@ II.2 Calculation
$\Pi^{(1a)}_{\mu\nu}(q^{2})=$ $\;{\mathrm{i}}e^{2}\int\dfrac{\mathrm{d}^{D}k}{(2\pi)^{D}}\left[-{\rm Tr}\left(\gamma_{\mu}[S(k)]^{2}\gamma_{\nu}S(k-q)\right)\right]\,,$ (10)
@@ II.3 Renormalization
$P^{(1a)}_{\rm R}(q^{2})=P^{(1a)}(q^{2})-P^{(1a)}(0)\\ =\frac{1-v^{2}}{4v}\biggl{\{}\frac{2v\left(3v^{2}-5\right)}{3\left(v^{2}-1\right)}+\left(v^{2}-1\right)\ln\left(\frac{v-1}{v+1}\right)\biggr{\}}\\ +\epsilon Q^{(1a)}_{1}(v)+\epsilon^{2}Q^{(1a)}_{2}(v)+{\mathcal{O}}(\epsilon^{3})\,.$ (24)
@@
$Q^{(1a)}_{1}(v)=\frac{1-v^{2}}{4v}\biggl{\{}\biggl{[}\left(v^{2}-1\right)\biggl{(}-2\text{Li}_{2}\left(\frac{1-v}{v+1}\right)\\ +\frac{1}{2}\ln^{2}\left(\frac{v-1}{v+1}\right)-\frac{\pi^{2}}{6}\biggr{)}+2\ln\left(\frac{v-1}{v+1}\right)\\ \times\biggl{(}v^{2}-\left(v^{2}-1\right)\ln\left(\frac{2v}{v+1}\right)\biggr{)}+4v\biggr{]}\biggr{\}}$ (25)
The quadratic term in $\epsilon$, denoted as $Q^{(2a)}_{1}(v)$, is given in Appendix A. Actually, only the terms proportional to $\epsilon$ are necessary for the renormalization of the two-loop vacuum polarization, but higher-order terms are useful for higher-loop calculations, and are thus relegated to Appendix A.
@@ III.1 Orientation
$\;\left[-{\rm Tr}\left(\gamma_{\mu}\,S(p_{4})\,\gamma^{\rho}\,S(p_{2})\,\gamma^{\sigma}\,S(p_{4})\,\gamma_{\nu}(-p_{3})\,D_{\rho\sigma}(p_{5})\right)\right]\,.$ (30b)
@@
$D_{\mu\nu}(k)=-\frac{1}{k^{2}+\mathrm{i}\epsilon}\left(g_{\rho\sigma}-\lambda\frac{k_{\rho}k_{\sigma}}{k^{2}}\right)\,,$ (31)
@@ III.3 Differential Equations and Master Integrals
A system of differential equations in $q^{2}$ is being created; we need to compute the 5 master integrals. Three of them are reducible into products of one-loop integrals, as outlined in Eq.(III.2. We obtain a system of differential equations in $q^{2}$ satisfied by the irreducible master integrals Ko1991plb ; Re1997ny ; GeRe2000 . The first of these equations is of second order,
@@ III.5 Renormalization
$P^{(2;1)}(0)+2P^{(2;2)}(0)=\frac{10\epsilon^{2}-7\epsilon-9}{12(\epsilon-2)\epsilon(2\epsilon+1)}\,.$ (46)
@@
$Z_{J}=1+\sum_{i}Z_{J}^{(L)}\left(\frac{\alpha}{\pi}\right)^{L}\left(\frac{e^{2}\,C(\epsilon)}{4\pi\alpha}\right)^{L}\,,$ (47)
@@ IV.1 Muonic Hydrogen
$\left.E^{({\rm(r)})}(2P-2S)\right|_{\mbox{$\mu$H}}=0.25495\,{\rm meV}\,.$ (75)
@@ IV.2 Relativistic Ordinary Hydrogen
$\Delta E^{{\rm(r)}}_{\rm NR}=\left(\dfrac{\alpha}{\pi}\right)^{2}\frac{(Z\alpha)^{4}}{n^{3}}\biggl{\{}(Z\alpha)B_{50}^{{\rm(r)}}+(Z\alpha)^{2}B_{60}^{(r)}\\ +(Z\alpha)^{3}\left[B_{71}^{(r)}\,\ln[(Z\alpha)^{-2}]+B_{70}^{(r)}\right]\biggr{\}}\,.$ (78)
@@
$B_{60}^{(r)}(nP_{1/2})=$ $\frac{8}{675}\frac{n^{2}-1}{n^{2}}\,,$ (80)
$B_{70}^{(r)}(nP_{1/2})=$ $-\frac{983\pi}{93555}\frac{n^{2}-1}{n^{2}}\,.$ (81)
@@
$B^{(2)}_{60}(nS)=$ $\;B_{40}(nS)\,W_{n}\ +\frac{77\zeta(3)}{72}-\frac{2311}{405}+\frac{1313}{675n^{2}}\,,$ (85c)
$B^{(2)}_{61}(nS)=$ $\;\tfrac{1}{2}B_{40}(nS)\,,\qquad B^{(2)}_{71}(nS)=\tfrac{1}{2}B_{50}(nS)\,.$ (85d)
@@
$B^{(2)}_{70}(nS)=$ $\;C_{50}\,W_{n}+C^{(a)}_{70}+\frac{C^{(b)}_{70}}{n^{2}}\,,$
$C^{(a)}_{70}=$ $\frac{104}{63}\pi\ln^{2}(2)-\frac{25}{63}\pi^{2}\ln(2)-\frac{200\,\pi}{63}\,G+\frac{13\pi^{3}}{189}$
@@
$C^{(b)}_{70}=$ $-\frac{3304831\pi}{1428840}+\frac{1891\pi^{2}}{2268}-\frac{1076}{567}\pi\ln(2)\,.$ (85f)
@@
$B^{(2)}_{70}(nP_{1/2})=$ $-\pi\biggl{[}\frac{1764797}{7144200}-\frac{809\pi}{11340}$
$+\frac{328\ln(2)}{2835}\biggr{]}\frac{n^{2}-1}{n^{2}}\,.$ (86b)
@@
$B^{(2)}_{70}(nP_{3/2})=$ $-\pi\left[\frac{75763}{510300}-\frac{31\pi}{810}+\frac{19\ln(2)}{405}\right]\frac{n^{2}-1}{n^{2}}\,.$ (87b)
@@ Acknowledgments
The authors acknowledge insightful conversations with Professor Ettore Remiddi. Support from the National Science Foundation (grant PHY–2110294) is gratefully acknowledged. S.L. also acknowledges support from Italian Ministry of University and Research (MUR) via the PRIN 2022 project n. 20225X52RA — MUS4GM2 funded by the European Union — Next Generation EU.
@@
$N_{12}(q^{2})$ $=-12\text{Li}_{3}\left(\frac{1-v}{v+1}\right)-24\text{Li}_{3}\left(\frac{2v}{v+1}\right)+\ln^{3}\left(\frac{v-1}{v+1}\right)-6\ln\left(\frac{2v}{v+1}\right)\ln^{2}\left(\frac{v-1}{v+1}\right)$
@@
For the term $Q_{1a}(r)$ from the integral with two fermion propagators, one finds for the term quadratic in $\epsilon$ the expression
$Q^{(1a)}_{2}(v)=$ $\;\frac{1-v^{2}}{4v}\,\biggl{\{}-\frac{1}{6}\left(v^{2}-1\right)\biggl{[}12\text{Li}_{3}\left(\frac{1-v}{v+1}\right)+24\text{Li}_{3}\left(\frac{2v}{v+1}\right)-\ln^{3}\left(\frac{v-1}{v+1}\right)+6\ln\left(\frac{2v}{v+1}\right)\biggl{(}\ln^{2}\left(\frac{v-1}{v+1}\right)-\pi^{2}\biggr{)}$
@@ Appendix B Coefficient Functions
$\left.+128q^{2}+176\right)+16\left(q^{8}-4q^{6}+6q^{4}+6q^{2}+60\right)\biggr{]}\,.$ (102)
@@ D.1 One–Loop Diagram
$E_{1}(2P-2S)=\langle 2P|V^{(1)}_{R}|2P\rangle-\langle 2S|V^{(1)}_{R}|2S\rangle\\ =\frac{(Z\alpha)^{2}\mu}{2\pi}\int\limits_{4m^{4}}^{\infty}\frac{\mathrm{d}(q^{2})}{q^{2}}\;\mathrm{Im}\!\left[\Pi_{\rm R}^{(1)}(q^{2}+\mathrm{i}\epsilon)\right]\,\frac{(\beta q)^{2}}{(1+\beta q)^{4}}\\ =\left(\dfrac{\alpha}{\pi}\right)\,(Z\alpha)^{2}\mu f_{1}(\beta)\,,\\ f_{1}(\beta)=\int_{4}^{{\infty}}\frac{\mathrm{d}(q^{2})}{q^{2}}\frac{{\rm Im}\Pi_{\rm R}^{(1)}(q^{2}+\mathrm{i}\epsilon)}{\alpha}\frac{(\beta q)^{2}}{2(1+\beta q)^{4}}\,.$ (111)
@@ D.4 Individual Energy Shifts
$\biggl{[}f^{(2)}_{2S}(\beta)=\frac{1}{1152w^{6}}\biggl{[}3\biggl{(}\biggl{(}653w^{10}-1058w^{8}+93w^{6}\\ +56w^{4}+256w^{2}-192\biggr{)}\left(\ln^{2}\left(\frac{1-w}{1+w}\right)+\pi^{2}\right)\\ -4w^{2}\left(447w^{6}-430w^{4}-88w^{2}+48\right)\biggr{)}\\ -384\left(55w^{6}-97w^{4}+28w^{2}+6\right)w^{3}I_{2}\left(\sqrt{\frac{1-w}{1+w}}\right)\\ +96\left(55w^{4}-110w^{2}+51\right)w^{6}I_{1}\left(\sqrt{\frac{1-w}{1+w}}\right)\\ +12\left(-227w^{8}+77w^{6}+224w^{4}+52w^{2}-120\right)w\times\\ \ln\left(\frac{1-w}{1+w}\right)+\sqrt{1-w^{2}}\biggl{(}\pi\biggl{(}2048w^{6}\ln(2)\\ +\frac{32}{3}\left(495w^{4}-757w^{2}+54\right)w^{4}\biggr{)}-16\pi^{2}\\ \times\left(165w^{8}-220w^{6}+21w^{4}+30w^{2}-36\right)\biggr{)}\biggr{]}\,.$ (130)

### v2

@@ II.2 Calculation
$\Pi^{(1a)}_{\mu\nu}(q^{2})=$ $\;{\mathrm{i}}e^{2}\int\dfrac{\mathrm{d}^{D}k}{(2\pi)^{D}}\left[-{\rm Tr}\left(\gamma_{\mu}[S(k)]^{2}\gamma_{\nu}S(k-q)\right)\right]\,.$ (10)
@@ II.3 Renormalization
$P^{(1a)}_{\rm R}(q^{2})=P^{(1a)}(q^{2})-P^{(1a)}(0)\\ =\frac{v^{2}-1}{4v}\biggl{\{}\frac{2v\left(3v^{2}-5\right)}{3\left(v^{2}-1\right)}+\left(v^{2}-1\right)\ln\left(\frac{v-1}{v+1}\right)\biggr{\}}\\ +\epsilon Q^{(1a)}_{1}(v)+\epsilon^{2}Q^{(1a)}_{2}(v)+{\mathcal{O}}(\epsilon^{3})\,.$ (24)
@@
$Q^{(1a)}_{1}(v)=\frac{v^{2}-1}{4v}\biggl{[}\left(v^{2}-1\right)\biggl{(}-2\text{Li}_{2}\left(\frac{1-v}{v+1}\right)\\ +\frac{1}{2}\ln^{2}\left(\frac{v-1}{v+1}\right)-\frac{\pi^{2}}{6}\biggr{)}+2\ln\left(\frac{v-1}{v+1}\right)\\ \times\biggl{(}v^{2}-\left(v^{2}-1\right)\ln\left(\frac{2v}{v+1}\right)\biggr{)}+4v\biggr{]}$ (25)
The quadratic term in $\epsilon$, denoted as $Q^{(1a)}_{2}(v)$, is given in Appendix A. Actually, only the terms proportional to $\epsilon$ are necessary for the renormalization of the two-loop vacuum polarization, but higher-order terms are useful for higher-loop calculations, and are thus relegated to Appendix A.
@@ III.1 Orientation
$\;\left[-{\rm Tr}\left(\gamma_{\mu}\,S(p_{4})\,\gamma^{\rho}\,S(p_{2})\,\gamma^{\sigma}\,S(p_{4})\,\gamma_{\nu}\,S(-p_{3})\,D_{\rho\sigma}(p_{5})\right)\right]\,.$ (30b)
@@
$D_{\rho\sigma}(k)=-\frac{1}{k^{2}+\mathrm{i}\epsilon}\left(g_{\rho\sigma}-\lambda\frac{k_{\rho}k_{\sigma}}{k^{2}}\right)\,,$ (31)
@@ III.3 Differential Equations and Master Integrals
A system of differential equations in $q^{2}$ is being created; we need to compute the 5 master integrals. Three of them are reducible into products of one-loop integrals, as outlined in Eq.(34. We obtain a system of differential equations in $q^{2}$ satisfied by the irreducible master integrals Ko1991plb ; Re1997ny ; GeRe2000 . The first of these equations is of second order,
@@ III.5 Renormalization
$P^{(2;1)}(0)+2P^{(2;2)}(0)=-\frac{10\epsilon^{2}-7\epsilon-9}{12(\epsilon-2)\epsilon(2\epsilon+1)}\,.$ (46)
@@
$Z_{J}=1+\sum_{L}Z_{J}^{(L)}\left(\frac{\alpha}{\pi}\right)^{L}\left(\frac{e^{2}\,C(\epsilon)}{4\pi\alpha}\right)^{L}\,,$ (47)
@@ IV.1 Muonic Hydrogen
$\left.E^{{\rm(r)}}(2P-2S)\right|_{\mbox{$\mu$H}}=0.25495\,{\rm meV}\,.$ (75)
@@ IV.2 Relativistic Ordinary Hydrogen
$\Delta E^{{\rm(r)}}=\left(\dfrac{\alpha}{\pi}\right)^{2}\frac{(Z\alpha)^{4}}{n^{3}}\biggl{\{}(Z\alpha)B_{50}^{{\rm(r)}}+(Z\alpha)^{2}B_{60}^{{\rm(r)}}\\ +(Z\alpha)^{3}\left[B_{71}^{{\rm(r)}}\,\ln[(Z\alpha)^{-2}]+B_{70}^{{\rm(r)}}\right]\biggr{\}}\,.$ (78)
@@
$B_{60}^{{\rm(r)}}(nP_{1/2})=$ $\frac{8}{675}\frac{n^{2}-1}{n^{2}}\,,$ (80)
$B_{70}^{{\rm(r)}}(nP_{1/2})=$ $-\frac{983\pi}{93555}\frac{n^{2}-1}{n^{2}}\,.$ (81)
@@
$B^{(2)}_{60}(nS)=$ $\;B^{(2)}_{40}(nS)\,W_{n}\ +\frac{77\zeta(3)}{72}-\frac{2311}{405}+\frac{1313}{675n^{2}}\,,$ (85c)
$B^{(2)}_{61}(nS)=$ $\;\tfrac{1}{2}B^{(2)}_{40}(nS)\,,\qquad B^{(2)}_{71}(nS)=\tfrac{1}{2}B^{(2)}_{50}(nS)\,.$ (85d)
@@
$B^{(2)}_{70}(nS)=$ $\;B^{(2)}_{50}\,W_{n}+B^{(a)}_{70}+\frac{B^{(b)}_{70}}{n^{2}}\,,$
$B^{(a)}_{70}=$ $\frac{104}{63}\pi\ln^{2}(2)-\frac{25}{63}\pi^{2}\ln(2)-\frac{200\,\pi}{63}\,G+\frac{13\pi^{3}}{189}$
@@
$B^{(b)}_{70}=$ $-\frac{3304831\pi}{1428840}+\frac{1891\pi^{2}}{2268}-\frac{1076}{567}\pi\ln(2)\,.$ (85f)
@@
$B^{(2)}_{70}(nP_{1/2})=$ $\pi\biggl{[}\frac{1764797}{7144200}-\frac{809\pi}{11340}+\frac{328\ln(2)}{2835}\biggr{]}\frac{n^{2}-1}{n^{2}}\,.$ (86b)
@@
$B^{(2)}_{70}(nP_{3/2})=$ $\pi\left[\frac{75763}{510300}-\frac{31\pi}{810}+\frac{19\ln(2)}{405}\right]\frac{n^{2}-1}{n^{2}}\,.$ (87b)
@@ Acknowledgments
The authors acknowledge insightful conversations with Professor Ettore Remiddi. Support from the National Science Foundation (grant PHY–2110294) is gratefully acknowledged. S.L. also acknowledges support from Italian Ministry of University and Research (MUR) via the PRIN 2022 project n. 20225X52RA — MUS4GM2 funded by the European Union via the Next Generation EU package.
@@
$N_{12}(v)$ $=-12\text{Li}_{3}\left(\frac{1-v}{v+1}\right)-24\text{Li}_{3}\left(\frac{2v}{v+1}\right)+\ln^{3}\left(\frac{v-1}{v+1}\right)-6\ln\left(\frac{2v}{v+1}\right)\ln^{2}\left(\frac{v-1}{v+1}\right)$
@@
For the term $Q_{2}^{(1a)}(v)$ from the integral with two fermion propagators, one finds for the term quadratic in $\epsilon$ the expression
$Q^{(1a)}_{2}(v)=$ $\;\frac{v^{2}-1}{4v}\,\biggl{\{}-\frac{1}{6}\left(v^{2}-1\right)\biggl{[}12\text{Li}_{3}\left(\frac{1-v}{v+1}\right)+24\text{Li}_{3}\left(\frac{2v}{v+1}\right)-\ln^{3}\left(\frac{v-1}{v+1}\right)+6\ln\left(\frac{2v}{v+1}\right)\biggl{(}\ln^{2}\left(\frac{v-1}{v+1}\right)-\pi^{2}\biggr{)}$
@@ Appendix B Coefficient Functions
$\left.+128q^{2}+176\right)+16\left(q^{8}-4q^{6}+6q^{4}+6q^{2}+60\right)\biggr{]}\,,$ (102)
@@ D.1 One–Loop Diagram
$E_{1}(2P-2S)=\langle 2P|V^{(1)}_{R}|2P\rangle-\langle 2S|V^{(1)}_{R}|2S\rangle\\ =\frac{(Z\alpha)^{2}\mu}{2\pi}\int\limits_{4}^{\infty}\frac{\mathrm{d}(q^{2})}{q^{2}}\;\mathrm{Im}\!\left[\Pi_{\rm R}^{(1)}(q^{2}+\mathrm{i}\epsilon)\right]\,\frac{(\beta q)^{2}}{(1+\beta q)^{4}}\\ =\left(\dfrac{\alpha}{\pi}\right)\,(Z\alpha)^{2}\mu f_{1}(\beta)\,,\\ f_{1}(\beta)=\int_{4}^{{\infty}}\frac{\mathrm{d}(q^{2})}{q^{2}}\frac{{\rm Im}\Pi_{\rm R}^{(1)}(q^{2}+\mathrm{i}\epsilon)}{\alpha}\frac{(\beta q)^{2}}{2(1+\beta q)^{4}}\,.$ (111)
@@ D.4 Individual Energy Shifts
$f^{(2)}_{2S}(\beta)=\frac{1}{1152w^{6}}\biggl{[}3\biggl{(}\biggl{(}653w^{10}-1058w^{8}+93w^{6}\\ +56w^{4}+256w^{2}-192\biggr{)}\left(\ln^{2}\left(\frac{1-w}{1+w}\right)+\pi^{2}\right)\\ -4w^{2}\left(447w^{6}-430w^{4}-88w^{2}+48\right)\biggr{)}\\ -384\left(55w^{6}-97w^{4}+28w^{2}+6\right)w^{3}I_{2}\left(\sqrt{\frac{1-w}{1+w}}\right)\\ +96\left(55w^{4}-110w^{2}+51\right)w^{6}I_{1}\left(\sqrt{\frac{1-w}{1+w}}\right)\\ +12\left(-227w^{8}+77w^{6}+224w^{4}+52w^{2}-120\right)w\times\\ \ln\left(\frac{1-w}{1+w}\right)+\sqrt{1-w^{2}}\biggl{(}\pi\biggl{(}2048w^{6}\ln(2)\\ +\frac{32}{3}\left(495w^{4}-757w^{2}+54\right)w^{4}\biggr{)}-16\pi^{2}\\ \times\left(165w^{8}-220w^{6}+21w^{4}+30w^{2}-36\right)\biggr{)}\biggr{]}\,.$ (130)